7x - 2 = 2y
3x = 2y - 1
x = ?
y = ?



Answer by checkley77(12844)

7x-2=2y
3x+1=2y SUBTRACT.
----------------
4X-3=0
4X=3
X=3/4 ANS.
7*3/4-2=2Y
21/4-2=2Y
2Y=21/4-8/4
2Y=13/4
Y=(13/4)/2
Y=13/4*1/2
Y=13/8 ANS.
PROOF:
3*3/4+1=2*13/8
9/4+1=26/8
9/4+4/4=13/4
13/4=13/4